### v3

Abstract
(I) We revisit the algorithmic problem of finding all triangles in a graph $G=(V,E)$ with $n$ vertices and $m$ edges. According to a result of Chiba and Nishizeki (1985), this task can be achieved by a combinatorial algorithm running in $O(m\alpha)=O(m^{3/2})$ time, where $\alpha=\alpha(G)$ is the graph arboricity. We provide a new very simple combinatorial algorithm for finding all triangles in a graph and show that is amenable to the same running time analysis. We derive these worst-case bounds from first principles and with very simple proofs that do not rely on classic results due to Nash-Williams from the 1960s. Our experimental results show that our simple algorithm for triangle listing is substantially faster in practice than that of Chiba and Nishizeki on all examples of real-world graphs we tried.
(II) We extend our arguments to the problem of finding all small complete subgraphs of a given fixed size. We show that the dependency on $m$ and $\alpha$ in the running time $O(\alpha^{\ell-2}\cdot m)$ of the algorithm of Chiba and Nishizeki for listing all copies of $K_{\ell}$, where $\ell\geq 3$, is asymptotically tight.
(III) We give improved arboricity-sensitive running times for counting and/or detection of copies of $K_{\ell}$, for small $\ell\geq 4$. A key ingredient in our algorithms is, once again, the algorithm of Chiba and Nishizeki. Our new algorithms are faster than all previous algorithms in certain high-range arboricity intervals for every $\ell\geq 7$.
Keywords: triangle, subgraph detection/counting, graph arboricity, rectangular matrix multiplication.
1 Introduction
The problem of deciding whether a given graph $G=(V,E)$ with $n$ vertices and $m$ edges contains a complete subgraph on $k$ vertices is among the most natural and easily stated algorithmic graph problems. If the subgraph size $k$ is part of the input, this is the Clique problem which is NP-complete [17]. For every fixed $k$, determining whether a given graph $G=(V,E)$ contains a complete subgraph on $k$ vertices can be accomplished by a brute-force algorithm running in $O(n^{k})$ time.
For $k=3$, deciding whether a graph contains a triangle and finding one if it does, or counting all triangles in a graph, can be done in $O(n^{\omega})$ time by the algorithm of Itai and Rodeh [16], where $\omega$ is the exponent of matrix multiplication [1, 7], and $\omega<2.372$ is the best known bound [10, 38]. The algorithm compares $M$ and $N=M^{2}$, where $M$ is the graph adjacency matrix; if there is a pair of indexes $i,j$ such that $M[i,j]=N[i,j]=1$, then there is a triangle based on edge $ij$. An equivalent characterization is: $G$ contains a triangle iff $M^{3}$ has a non-zero entry on its main diagonal. See [27, Ch. 10] for a short exposition of this elegant method. Alternatively, this task can be done in $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time by the algorithm of Alon, Yuster, and Zwick [3], which deals with vertices of high and low degree separately. Itai and Rodeh [16] and also Papadimitriou and Yannakakis [32] as well as Chiba and Nishizeki [6] showed that triangles in planar graphs can be found in $O(n)$ time.
For $k=4$, deciding whether a graph contains a $K_{4}$ and finding one if it does (or counting all $K_{4}$’s in a graph) can be done in $O(n^{\omega+1})=O(n^{3.373})$ time by the algorithm of Alon, Yuster, and Zwick [3], in $O(n^{3.252})$ time by the algorithm of Eisenbrand and Grandoni [12], and in $O(m^{(\omega+1)/2})=O(m^{1.686})$ time by the algorithm of Kloks, Kratsch, and Müller [18].
In contrast to the problem of detecting the existence of subgraphs of a certain kind, the analogous problem of listing all such subgraphs has usually higher complexity, as expected. For example, finding all triangles in a given graph (each triangle appears in the output list) can be accomplished in $O(m^{3/2})$ time and with $O(n^{2})$ space by an extended version of the algorithm of Itai and Rodeh [16]. Bar-Yehuda and Even [4] improved the space complexity of the algorithm from $O(n^{2})$ to $O(n)$ by avoiding the use of the adjacency matrix.
Chiba and Nishizeki [6] further refined the time complexity in terms of graph arboricity, which is the minimum number of edge-disjoint forests into which its edges can be partitioned. Their algorithm lists all triangles in a graph in $O(m\alpha)$ time, where $\alpha$ is the arboricity of the graph. They also showed that $\alpha=O(\sqrt{m})$; consequently, the running time is $O(m^{3/2})$. Since there are graphs $G$ with $\alpha(G)=\Theta(m^{1/2})$, this does not improve the worst-case dependence on $m$. Moreover, since there are graphs with $\Theta(m^{3/2})$ triangles, see, e.g., [33], this dependence cannot be improved. For every fixed $\ell\geq 3$, the same authors gave an algorithm for listing all copies of $K_{\ell}$ in $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$ time. If $G$ is planar, $\alpha(G)\leq 3$ (see [15, p. 124]), so their algorithm runs in $O(m)=O(n)$ (i.e., linear) time on planar graphs.
We distinguish several variants of the general problem of finding triangles in a given undirected graph $G=(V,E)$: (i) the triangle detection problem is that of deciding if $G$ is triangle-free (resp., finding a triangle in $G$ otherwise); (ii) the triangle counting problem is that of determining the total number of triangles in $G$; (iii) the triangle listing problem is that of listing (finding, reporting) all triangles in $G$, with each triangle appearing in the output list. Obviously, any algorithm for listing all triangles can be easily transformed into one for triangle detection or into one for listing a specified number of triangles, by stopping after the required number of triangles have been output.
Our results.
We obtain the following results for the problem of finding small complete subgraphs of a given size. Each of our algorithms for deciding whether a graph contains a given complete subgraph (item (iii) below) also finds a suitable witness, if it exists.
  1. (i)
    We provide a new combinatorial algorithm for finding all triangles in a graph running in $O(m\alpha)=O(m^{3/2})$ time, where $\alpha=\alpha(G)$ is the graph arboricity (Algorithm Hybrid in Section 2). We derive these worst-case bounds from first principles and with very simple proofs that do not rely on classic results due to Nash-Williams from the 1960s. We present also an experimental comparison of our new simple algorithm for finding all triangles with that of Chiba and Nishizeki. It shows that our algorithm is substantially faster in practice on all considered examples of real-world graphs.
  2. (ii)
    For every $n$, $b\leq n/2$ and a fixed $\ell\geq 3$, there exists a graph $G$ of order $n$ with $m$ edges and $\alpha(G)\leq b$ that has $\Omega(\alpha^{\ell-2}\cdot m)$ copies of $K_{\ell}$ (Lemma 3 in Section 3). As such, the dependency on $m$ and $\alpha$, in the running time $O(\alpha^{\ell-2}\cdot m)$ for listing all copies of $K_{\ell}$, is asymptotically tight.
  3. (iii)
    We give improved arboricity-sensitive running times for counting and/or detection of copies of $K_{\ell}$, for small $\ell\geq 4$ (Section 4). A key ingredient in our algorithms is the algorithm of Chiba and Nishizeki. Our new algorithms beat all previous algorithms in certain high-range arboricity intervals for every $\ell\geq 7$. We also provide up-to-date running times based on rectangular matrix multiplication times.
Preliminaries and notation.
Let $G=(V,E)$ be an undirected graph. The neighborhood of a vertex $v\in V$ is the set $N(v)=\{w\ :\ (v,w)\in E\}$ of all adjacent vertices, and its cardinality $\deg(v)=|N(v)|$ is the degree of $v$ in $G$. A clique in a graph $G=(V,E)$ is a subset $C\subset V$ of vertices, each pair of which is connected by an edge in $E$. The Clique problem is to find a clique of maximum size in $G$. An independent set of a graph $G=(V,E)$ is a subset $I\subset V$ of vertices such that no two of them are adjacent in $G$.
Unless specified otherwise: (i) all algorithms discussed are assumed to be deterministic; (ii) all graphs mentioned are undirected; (iii) all logarithms are in base $2$.
It should be noted that the asymptotically fastest algorithms for matrix multiplication have mainly a theoretical importance and are otherwise largely impractical [26, Ch. 10.2.4]. As such, any algorithm that employs fast matrix multiplication (FMM) inherits this undesirable feature. Whereas the notion of “combinatorial” algorithm does not have a formal or precise definition, such an algorithm is usually simpler to implement and practically efficient, see [37]. Combinatorial algorithms are sometimes theoretically less efficient than non-combinatorial ones but practically more efficient since they tend to be simpler to implement.
Graph parameters.
For a graph $G$, its arboricity $\alpha(G)$ is the minimum number of edge-disjoint forests into which $G$ can be decomposed [6]. For instance, it is known and easy to show that $\alpha(K_{n})=\lceil n/2\rceil$. A characterization of arboricity is provided by the following classic result [28, 29, 36]; see also [9, p. 60].
Theorem 1.
(Nash-Williams 1964; Tutte 1961) A multigraph $G=(V,E)$ can be partitioned into at most $k$ forests if and only if every set $U\subseteq V$ induces at most $k(|U|-1)$ edges.
The degeneracy $d(G)$ of an undirected graph $G$ is the smallest number $d$ for which there exists an acyclic orientation of $G$ in which all the out-degrees are at most $d$. The degeneracy of a graph is linearly related to its arboricity, i.e., $\alpha(G)=\Theta(d(G))$; more precisely $\alpha(G)\leq d(G)\leq 2\alpha(G)-1$; see [3, 13, 25, 41].
It is worth noting that high arboricity is not an indication for the graph having many (or any) triangles. Indeed, the complete bipartite graph $K_{\lfloor n/2\rfloor,\lceil n/2\rceil}$, or any of its dense subgraphs has arboricity $\Theta(n)$ but no triangles.
1.1 Triangle Finding Algorithms
To place our algorithm in Section 2 in a proper context, we first present a summary of previous work in the area of triangle finding and enumeration.
The algorithms of Itai and Rodeh (1978).
The authors gave three methods for finding triangles. Initially intended for triangle detection, the first algorithm [16] runs in $O(m^{3/2})$ time. It can be extended to list all triangles within the same overall time and works as follows:
Itai-Rodeh$(G)$
Input: an undirected graph $G=(V,E)$
1 Find a spanning tree for each connected component of $G$
2 List all triangles containing at least one tree edge
3 Delete the tree edges from $G$ and go to Step 1
The second is a randomized algorithm that checks whether there is an edge contained in a triangle. It runs in $O(mn)$ worst-case time and $O(n^{5/3})$ expected time. The third algorithm relies on Boolean matrix multiplication and runs in $O(n^{\omega})$ time. (Boolean matrix multiplication can be reduced to standard matrix multiplication; see [26, Ch. 10.2.4] for details.)
The algorithm of Chiba and Nishizeki (1985).
The algorithm uses a vertex-iterator approach for listing all triangles in $G$. It relies on the observation that each triangle containing a vertex $v$ corresponds to an edge joining two neighbors of $v$. The graph is represented with doubly-linked adjacency lists and mutual references between the two stubs of an edge ensure that each deletion takes constant time. A more compact version described by Ortmann and Brandes [31] is given below.
Chiba-Nishizeki$(G)$
Input: an undirected graph $G=(V,E)$
1 Sort vertices such that $\deg(v_{1})\geq\deg(v_{2})\geq\ldots\deg(v_{n})$
2 for $u=v_{1},v_{2},\ldots,v_{n-2}$ do
3      foreach vertex $v\in N(u)$ do mark $v$
4      foreach vertex $v\in N(u)$ do
5           foreach vertex $w\in N(v)$ do
6                if $w$ is marked then output triangle $uvw$
7
8                end foreach
9               unmark $v$
10                end foreach
11               $G\leftarrow G-u$
12                end for
The authors [6] showed that their algorithm runs in $O(m\alpha)$ time. As a corollary, the number of triangles is $O(m\alpha)$ as well. The $O(m^{3/2})$ upper bound on the number of triangles in a graph is likely older than these references indicate. In any case, other proofs are worth mentioning [11, 19], including algebraic ones [33]. We derive yet another one in Section 2.
Corollary 1.
([16, 6]) For any graph $G$ of order $n$ with $m$ edges and arboricity $\alpha$, $G$ contains $O(m\alpha)=O(m^{3/2})$ triangles.
Ortmann and Brandes [31] gave a survey of other approaches, including edge-iterators. Algorithms in this category iterate over all edges and intersect the neighborhoods of the endpoints of each edge. A straightforward neighborhood merge requires $O(\deg(u)+\deg(v))$ time per edge $uv$, but this is not good enough to list all triangles in $O(m^{3/2})$ time. Two variants developed by Shanks [34] use $O(m)$ extra space to represent neighborhoods in hash sets and obtain the intersection in $O(\min(\deg(u),\deg(v)))$ time, which suffices for listing all triangles in $O(m^{3/2})$ time.
The algorithm of Alon, Yuster and Zwick (1997).
The authors showed that deciding whether a graph contains a triangle and finding one if it does (or counting all triangles in a graph) can be done in $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time [3]. The idea is to find separately triangles for which at least one vertex has low degree (for an appropriately set threshold) and triangles whose all three vertices have high degree. Triangles of the latter type are handled by applying matrix multiplication to a smaller subgraph.
Recent algorithms.
Björklund et al. [5] obtained output-sensitive algorithms for finding (and listing) all triangles in a graph; their algorithms are tailored for dense and sparse graphs. Several approaches [14, 20] provide asymptotic improvements by taking advantage of the bit-level parallelism offered by the word-RAM model. Although they do not appear to be very practical, these methods asymptotically improve on the $O(m\alpha+k)$ running time of the earlier algorithms (in particular that in [6]). If the number of triangles is small, Zechner and Lingas [40] showed how to list all triangles in $O(n^{\omega})$ time.
2 A Simple Hybrid Algorithm for Listing All Triangles
In this section we present a new algorithm for listing all triangles. While its general idea is not new, the specific hybrid data structure therein does not appear to have been previously considered. Using both an adjacency list representation and an adjacency matrix representation of the graph yields a very time-efficient neighborhood merge (intersection) procedure. Let $V=\{1,2,\ldots,n\}$ and let $M$ be the adjacency matrix of $G$. A triangle $ijk$ with $i<j<k$ is reported when edge $ij$ is processed, and so each triangle is reported exactly once. For each edge $ij$, the algorithm scans the adjacency list of the endpoint of lower degree (among $i$ and $j$) and for each neighbor it checks for the needed entry in the adjacency matrix $M$ corresponding to the third triangle edge. In the pseudocode below, $ADJ(v)$ denotes the adjacency list of vertex $v$.
Hybrid$(G)$
Input: an undirected graph $G=(V,E)$
1 foreach edge $ij\in E$, ($i<j$do
2      if $\deg(i)\leq\deg(j)$ then $x\leftarrow i$, $y\leftarrow j$
3      else $x\leftarrow j$, $y\leftarrow i$
4      foreach $k\in ADJ(x)$ do
5           if $j<k$ and $M(y,k)=1$ then report triangle $ijk$
6
7           end foreach
8
9           end foreach
We subsequently show that the algorithm runs in time $O(m\alpha)$. Whereas the space used is quadratic, the hybrid algorithm appears to win by running time and simplicity. In particular, no additional data structures nor hashing are used, no sorting (by degree or otherwise) and no doubly linked lists are needed, etc. Similar algorithms that use hashing in order to run in linear space are $6$ to $8$ times slower than their “unoptimized” counterparts on most graphs with about $5\times 10^{6}$ vertices in the study in [34, 35]. “The two algorithms edge-iterator-hashed and forward-hashed, which use hashed data structures for every node, perform badly” on such graphs, the authors write [35, p. 8].
2.1 The Analysis
Define the following function $f$ on the edges of $G$. For $uv=e\in E$, let
$f(e)=\min(\deg(u),\deg(v))\text{ and }F(G)=\sum_{e\in E}f(e).$ (1)
There are at most $f(e)$ triangles based on edge $e=ij$ and overall at most $F(G)$ triangles in $G$. The runtime of the algorithm is proportional to this quantity; the space used is quadratic. A short and elegant decomposition argument by Chiba and Nishizeki [6, Lemma 2] shows that
$F(G)\leq 2m\alpha,$ (2)
thus our algorithm runs in time $O(m\alpha)$. The same analysis applies to the algorithm of Chiba and Nishizeki [6]. By Theorem 1, the above authors deduced that $\alpha(G)\leq\lceil(2m+n)^{1/2}/2\rceil$, which implies that $\alpha(G)=O(m^{1/2})$ for a connected graph $G$. As such, both algorithms run in $O(m^{3/2})$ time on any graph.
Lemma 1 below shows how to bypass the Nash-Williams arboricity bound (Theorem 1) and deduce the $O(m^{3/2})$ upper bound for listing all triangles in a graph from first principles.
Lemma 1.
Let $G$ be a graph on $n$ vertices with $m$ edges. Then
$F(G)\leq 4m^{3/2}.$
Proof.
(folklore) There are two types of edges $uv$:
  1. 1.
    $\min(\deg(u),\deg(v))\leq 2\sqrt{m}$
  2. 2.
    $\min(\deg(u),\deg(v))>2\sqrt{m}$
There are at most $m$ edges of type 1 and each contributes at most $2\sqrt{m}$ to $F(G)$, so the total contribution from edges of this type is at most $2m^{3/2}$.
Each edge of type 2 connects two nodes of degree $>2\sqrt{m}$, and there are at most $2m/(2\sqrt{m})=\sqrt{m}$ such nodes. The degree of each of them thus contributes to $F(G)$ at most $\sqrt{m}$ times and the sum of all degrees of such nodes is at most $2m$. It follows that the total contribution from edges of type 2 is at most $2m^{3/2}$.
Overall, we have $F(G)\leq 4m^{3/2}$. ∎
Remarks.
Lemma 1 immediately gives an upper bound of $4m^{3/2}$ on the number of triangles in a graph. It is in fact known [33] that this number is at most $(2^{1/2}/3)\,m^{3/2}$ (and this bound is sharp). More generally, for every fixed $\ell\geq 3$, the number of copies of $K_{\ell}$ is $O(\alpha^{\ell-2}\cdot m)=O(m^{\ell/2})$; see [6, Thm. 3].
2.2 Experimental Results
Although Hybrid and Chiba-Nishizeki have the same asymptotic running time, we claim that our algorithm is faster in practice due to its simplicity. To support this claim, we implemented both algorithms in C++ and compared their running time on complete graphs $K_{1000}$ and $K_{2000}$, as well as a collection of real-world graphs with various sizes. The code is publicly available at https://gitlab.com/kanatos92/triangle-listing. The graphs are taken from the SNAP datasets [23] with self-loops and multi-edges removed. Graph statistics are listed in Table 1.
Graph Vertices Edges Triangles
$K_{1000}$ 1,000 499,500 166,167,000
$K_{2000}$ 2,000 1,999,000 1,331,334,000
facebook-combined 4,039 88,234 1,612,010
musae-facebook 22,470 170,823 794,953
cit-HepPh 34,546 420,877 1,276,868
large-twitch 168,114 6,797,557 54,148,895
loc-gowalla 196,591 950,327 2,273,138
com-dblp 317,080 1,049,866 2,224,385
higgs-social-network 456,626 12,508,413 83,023,401
web-BerkStan 685,230 6,649,470 64,690,980
com-youtube 1,134,890 2,987,624 3,056,386
as-skitter 1,696,415 11,095,298 28,769,868
roadNet-CA 1,965,206 2,766,607 120,676
Table 1: Statistics of the graphs.
For each algorithm, we measure the construction time taken to build the needed graph representation, and the listing time taken to enumerate all the triangles. To minimize the effect of system fluctuations, I/O operations are not timed. In particular, each input graph is preloaded into memory as a list of edges, and the triangles are stored in a container (instead of being sent directly to output) during enumeration. The experiments are conducted on an AWS server with an AMD EPYC 7R13 processor and 512 GB memory. For a fair comparison, both algorithms are run on a single processor. However, it is worth noting that the loop in Hybrid can easily be made to run in parallel (by adding a single line of pragma directive from OpenMP, for example) whereas it is not so trivial for Chiba-Nishizeki because the algorithm modifies its data structures in each iteration.
Both algorithms are repeated $10$ times on each graph, the average time is summarized in Table 2. Observe that, although for some graphs Hybrid takes more time in the construction phase, it is the overall winner for all the inputs in the table. In most cases, the running time of Hybrid is less than half of the running time of Chiba-Nishizeki.
Graph Chiba-Nishizeki Hybrid
Construction Listing Total Construction Listing Total
$K_{1000}$ $37.4$ $11221.8$ $11259.2$ $\mathbf{8.0}$ $\mathbf{5778.3}$ $\mathbf{5786.3}$
$K_{2000}$ $146.6$ $120570.2$ $120716.8$ $\mathbf{34.6}$ $\mathbf{46121.5}$ $\mathbf{46156.1}$
facebook-combined $6.0$ $89.9$ $95.9$ $\mathbf{1.7}$ $\mathbf{54.3}$ $\mathbf{56.0}$
musae-facebook $16.7$ $104.3$ $121.0$ $\mathbf{13.3}$ $\mathbf{44.5}$ $\mathbf{57.8}$
cit-HepPh $45.8$ $888.4$ $934.2$ $\mathbf{41.2}$ $\mathbf{202.1}$ $\mathbf{243.3}$
large-twitch $913.9$ $48107.8$ $49021.7$ $\mathbf{898.3}$ $\mathbf{18424.4}$ $\mathbf{19322.7}$
loc-gowalla $\mathbf{138.3}$ $2500.8$ $2639.1$ $150.5$ $\mathbf{951.4}$ $\mathbf{1101.9}$
com-dblp $\mathbf{173.4}$ $1189.5$ $1362.9$ $224.1$ $\mathbf{454.8}$ $\mathbf{678.9}$
higgs-social-network $2387.1$ $144523.7$ $146910.8$ $\mathbf{2363.5}$ $\mathbf{54425.8}$ $\mathbf{56789.3}$
web-BerkStan $1219.5$ $16841.5$ $18061.0$ $\mathbf{923.3}$ $\mathbf{5828.8}$ $\mathbf{6752.1}$
com-youtube $\mathbf{557.6}$ $7915.6$ $8473.2$ $617.7$ $\mathbf{3106.1}$ $\mathbf{3723.8}$
as-skitter $\mathbf{2532.2}$ $47561.0$ $50093.2$ $2577.4$ $\mathbf{17494.4}$ $\mathbf{20071.8}$
roadNet-CA $\mathbf{690.0}$ $1187.3$ $1877.3$ $1210.8$ $\mathbf{226.9}$ $\mathbf{1437.7}$
Table 2: Average running time of both algorithms in milliseconds. The faster ones in each row are highlighted in bold.
3 Constructions
Lemma 2.
For every $n\geq 3$ and $3\leq m\leq{n\choose 2}$, there exists a graph $G$ of order $n$ with $m$ edges that contains $\Theta(m^{3/2})$ triangles.
Proof.
Let $3\leq x\leq n$ be the unique integer such that (${x\choose 2}\leq m<{x+1\choose 2}$. Note that $x=\Theta(\sqrt{m})$. Let $G=(V,E)$ where $V=V_{1}\cup V_{2}$, $|V_{1}|=x$, and $|V_{2}|=n-x$. Set $V_{1}$ induces a complete subgraph and set $V_{2}$ induces an independent set in $G$. Add $m-{x\choose 2}$ edges to $G$ arbitrarily so that $G$ has $m$ edges. Let $T$ be the set of triangles in $G$. We have
$|T|\geq{x\choose 3}=\Omega(m^{3/2}).$
By Lemma 1, $G$ contains $\Theta(m^{3/2})$ triangles, as required. ∎
Lemma 3.
Let $\ell\geq 3$ be a fixed integer. For every $n$ and $b\leq n/2$, there exists a graph $G$ of order $n$ with $\alpha\leq b$ that has $\Omega(\alpha^{\ell-2}\cdot m)$ copies of $K_{\ell}$, where $m$ is the number of edges in $G$ and $\alpha$ is the arboricity of $G$.
Proof.
We may assume that $b$ is even and $b/2\ |\ n$. Suppose that $n=(2k+1)b/2$. Let $G=(V,E)$, where $V=V_{1}\cup V_{2}$, and $|V_{1}|=kb$, $|V_{2}|=b/2$. Set $V_{1}$ consists of $k$ complete subgraphs of order $b$ whereas $V_{2}$ is an independent set of size $b/2$ in $G$. All edges in $V_{1}\times V_{2}$ are present in $G$. See Fig. 1.
Figure 1: Illustration of the graph $G$ for $k=3$, $b=4$.
Let $K$ be the set of complete subgraphs of order $\ell$. We have
$m$ $=|E|=k{b\choose 2}+k\frac{b^{2}}{2}=\Theta(k\,b^{2}),$
$|K|$ $=k{b\choose\ell}+k{b\choose\ell-1}\frac{b}{2}=\Omega(k\,b^{\ell})=\Omega(b^{\ell-2}\,m),$
Note that the arboricity of $G$ is bounded from above by $b$. Indeed, $\alpha\leq\frac{b}{2}+\frac{b}{2}=b$, since $\alpha(K_{b})=b/2$ and the edges in $V_{1}\times V_{2}$ can be partitioned into $b/2$ stars centered at the vertices in $V_{2}$. ∎
4 Finding Small Complete Subgraphs Efficiently
In this section we address the problem of detecting the presence of $K_{\ell}$ for a fixed $\ell\geq 4$. We combine and refine several approaches existent in the literature of the last $40$ years to obtain faster algorithms in general and for a large class of graphs with high arboricity. In particular, we will use the algorithm of Chiba and Nishizeki [6] for listing all copies of $K_{\ell}$ in $O(\alpha^{\ell-2}\cdot m)$ time. Our algorithms are formulated for the purpose of counting but they can be easily adapted for the purpose of detection.
The basic idea is as follows: since listing is generally slower than counting or detection, one can use a hybrid—decomposition based—approach that combines listing in Part A with counting or detection in Part B.
Recall that $\omega<2.372$ is the exponent of matrix multiplication [1, 7, 38], namely the infimum of numbers $\tau$ such that two $n\times n$ real matrices can be multiplied in $O(n^{\tau})$ time (operations). Similarly, let $\omega(p,q,r)$ stand for the infimum of numbers $\tau$ such that an $n^{p}\times n^{q}$ matrix can be multiplied by an $n^{q}\times n^{r}$ matrix in $O(n^{\tau})$ time (operations). For simplicity and as customary (see, e.g., [3, 30]), we write that two $n\times n$ matrices can be multiplied in $O(n^{\omega})$ time, since this does not affect our results; and similarly when multiplying rectangular matrices.
The extension method.
The algorithm solves the detection (or counting) problem for a complete subgraph $K_{\ell}$. Let $T(n,m,\ell)$ denote the running time of the algorithm for a graph with $n$ vertices and $m$ edges. Write $\ell=\ell_{1}+\ell_{2}$, for a suitable choice $\ell_{1},\ell_{2}\geq 2$. At the beginning, we run the algorithm of Chiba and Nishizeki to form a list of subgraphs isomorphic to $K_{\ell_{1}}$. Then, for each subgraph $G_{1}=(V_{1},E_{1})$ on the list: (i) we construct the subgraph $G_{2}$ of $G$ induced by vertices in $V\setminus V_{1}$ that are adjacent to all vertices in $V_{1}$; and (ii) we count the subgraphs isomorphic to $K_{\ell_{2}}$ in $G_{2}$ (or find one such subgraph); this is a (possibly recursive) call of the same procedure on a smaller instance. In other words, we count the number of extensions of the subgraph isomorphic to $K_{\ell_{1}}$ to a $K_{\ell}$. Another formulation of this method can be found in [21].
There are $O(\alpha^{\ell_{1}-2}\cdot m)$ copies of $K_{\ell_{1}}$ and they can be found in $O(\alpha^{\ell_{1}-2}\cdot m)$ time. For each fixed copy of $K_{\ell_{1}}$, the time spent in $G_{2}$ is at most $T(n,m,\ell_{2})$, and so the overall time satisfies the recurrence
$T(n,m,\ell)=O(\alpha^{\ell_{1}-2}\cdot m\cdot T(n,m,\ell_{2})).$
Each copy of $K_{\ell}$ is generated exactly ${\ell\choose\ell_{1}}$ times and so the total count needs to be divided by this number in the end.
The triangle method and its refinement.
The algorithm solves the detection (or counting) problem for a complete subgraph $K_{\ell}$. Nešetřil and Poljak [30] showed an efficient reduction of detecting and counting copies of any complete subgraph to the aforementioned method of Itai and Rodeh [16] for triangle detection and counting. To start with, consider the detection of complete subgraphs of size $\ell=3j$. For a given graph $G$ with $n$ vertices, construct an auxiliary graph $H$ with $O(n^{j})$ vertices, where each vertex of $H$ is a complete subgraph of order $j$ in $G$. Two vertices $V_{1},V_{2}$ in $H$ are connected by an edge if $V_{1}\cap V_{2}=\emptyset$ and all edges in $V_{1}\times V_{2}$ are present in $G$; equivalently, $V_{1}\cup V_{2}$ is a complete subgraph on $|V_{1}|+|V_{2}|$ vertices. The detection (or counting) of triangles in $H$ yields an algorithm for the detection (or counting) of $K_{\ell}$’s in $G$, running in $O(n^{j\omega})$ time. For detecting complete subgraphs of size $\ell=3j+i$, where $i\in\{1,2\}$, the algorithm can be adapted so that it runs in $O(n^{j\omega+i})$ time.
The triangle method can be extended to larger subgraphs, i.e., if $\ell=kj+i$, for some $k\geq 4$, and $i\in\{0,1,\ldots,k-1\}$, the algorithm tries to detect a $K_{k}$ in $H$; see also [21]. We give a few applications at the end of this section.
Here we focus on $k=3$. For convenience, define the following integer functions
$\beta(\ell)$ $=\omega(\lfloor\ell/3\rfloor,\lceil(\ell-1)/3\rceil,\lceil\ell/3\rceil),$ (3)
$\gamma(\ell)$ $=\lfloor\ell/3\rfloor\omega+\ell\pmod{3}.$ (4)
With this notation, the algorithm runs in $O(n^{\gamma(\ell)})$ time. Two decades later, Eisenbrand and Grandoni [12] refined the triangle method by using fast algorithms for rectangular matrix multiplication instead of those for square matrix multiplication. It partitions the graph into three parts of roughly the same size: $\ell_{1}=\lfloor\ell/3\rfloor$, $\ell_{2}=\lceil(\ell-1)/3\rceil$, $\ell_{3}=\lceil\ell/3\rceil$. The refined algorithm runs in time $O(n^{\beta(\ell)})$ time. If the rectangular matrix multiplication is carried out via the straightforward partition into square blocks and fast square matrix multiplication (see, e.g., [8, Exercise 4.2-6]), one recovers the time complexity of the algorithm of Nešetřil and Poljak; that is, $\beta(\ell)\leq\gamma(\ell)$, see [12] for details. Eisenbrand and Grandoni [12] showed that the above inequality is strict for a certain range: if $\ell\pmod{3}=1$ and $\ell\leq 16$, or $\ell\pmod{3}=2$. In summary, their algorithm is faster than that of Nešetřil and Poljak in these cases. We subsequently refer to their method as the refined triangle method. Another extension of the triangle method is considered in [21].
A problem of Kloks, Kratsch, and Müller.
The authors asked [18] whether there is an $O(m^{\ell\omega/6})$ algorithm for deciding whether a graph contains a $K_{\ell}$, if $\ell$ is a multiple of $3$. Eisenbrand and Grandoni showed that this is true for every multiple of $3$ at least $6$. Indeed, by their Theorem 2 [12], this task can be accomplished in time $O(m^{\beta(\ell)/2})$ for every $\ell\geq 6$, with $\beta(\ell)$ as in (3). If $\ell=3j$, where $j\geq 2$, then
$\beta(\ell)=\omega(\lfloor\ell/3\rfloor,\lceil(\ell-1)/3\rceil,\lceil\ell/3\rceil)=\omega(j,j,j)=j\cdot\omega(1,1,1)=j\omega.$
It follows that the running time is $O(m^{\beta(\ell)/2})=O(m^{j\omega/2})=O(m^{\ell\omega/6})$. The proof of the theorem is rather involved. Here we provide an alternative simpler argument that also yields an arboricity-sensitive bound, item (i) below.
General case derivations.
We first consider the general cases: (i) $\ell=3j$, $j\geq 3$; (ii) $\ell=3j+1$, $j\geq 2$; and (iii) $\ell=3j+2$, $j\geq 2$. It will be evident that our algorithms provide improved bounds for every $\ell\geq 7$. For a given $j\geq 2$, we shall consider the interval
$I_{j}=\left(\frac{\omega-1}{j(3-\omega)+2(\omega-1)},\,\frac{1}{2}\right).$
  1. (i)
    $\ell=3j$, $j\geq 3$. We use the triangle method with a refined calculation. The vertices of the auxiliary graph $H$ are subgraphs isomorphic to $K_{j}$. By the result of Chiba and Nishizeki [6], $H$ has $O(\alpha^{j-2}\cdot m)$ vertices. The algorithm counts triangles in $H$ in time proportional to
    $\left(\alpha^{j-2}\cdot m\right)^{\omega}=\alpha^{(j-2)\omega}\cdot m^{\omega}.$
    For $\ell=9,12,15$ (the entries in Table 3), these runtimes are $O(\alpha^{\omega}\cdot m^{\omega})$, $O(\alpha^{2\omega}\cdot m^{\omega})$, and $O(\alpha^{3\omega}\cdot m^{\omega})$, respectively.
    Since $\alpha=O(m^{1/2})$, the above expression is bounded from above as follows:
    $\alpha^{(j-2)\omega}\cdot m^{\omega}=O\left(\left(m^{(j-2)/2}\cdot m\right)^{\omega}\right)=O\left(m^{j\omega/2}\right)=O\left(m^{\ell\omega/6}\right).$
    Next, we show that for a certain high-range of $\alpha$, the new bound $\alpha^{(j-2)\omega}\cdot m^{\omega}$ beats all previous bounds, namely $m^{j\omega/2}$ and $\alpha^{3j-2}\cdot m$, for $j\geq 3$ (or $\ell=9,12,15,\ldots$). Let $\alpha=\Theta(m^{x})$, where $x\in I_{j}$. We first verify that
    $\alpha^{(j-2)\omega}\cdot m^{\omega}\ll m^{j\omega/2},\text{ or }((j-2)x+1)\omega<j\omega/2,$
    which holds for $x<1/2$. Secondly, we verify that
    $\alpha^{(j-2)\omega}\cdot m^{\omega}$ $\ll\alpha^{3j-2}\cdot m,\text{ or }((j-2)x+1)\omega<(3j-2)x+1,\text{ or }$
    $x$ $>\frac{\omega-1}{j(3-\omega)+2(\omega-1)},$
    which holds by the choice of the interval $I_{j}$. In particular, for $\ell=9$, this occurs for $x\in\left(\frac{\omega-1}{7-\omega},\frac{1}{2}\right)\supset(0.297,0.5)$; for $\ell=12$, this occurs for $x\in\left(\frac{\omega-1}{10-2\omega},\frac{1}{2}\right)\supset(0.262,0.5)$; for $\ell=15$, this occurs for $x\in\left(\frac{\omega-1}{13-3\omega},\frac{1}{2}\right)\supset(0.234,0.5)$. Moreover, if $\omega=2$, as conjectured, these intervals extend to $(1/5,1/2)$, $(1/6,1/2)$, and $(1/7,1/2)$, respectively.
  2. (ii)
    $\ell=3j+1$, $j\geq 2$. The refined triangle method [12] with $\ell_{1}=j$, $\ell_{2}=j+1$, $\ell_{3}=j$, leads to rectangular matrix multiplication $[O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]\cdot[O(\alpha^{j-1}m)\times O(\alpha^{j-2}m)]$. Its complexity is at most $O(\alpha)$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
    $T(n,m,\ell)=O(\alpha\cdot\alpha^{(j-2)\omega}\cdot m^{\omega})=O(\alpha^{(j-2)\omega+1}\cdot m^{\omega}).$
    For $\ell=7,10,13,16$ (the entries in Table 3), these runtimes are $O(\alpha\cdot m^{\omega})$, $O(\alpha^{\omega+1}\cdot m^{\omega})$, $O(\alpha^{2\omega+1}\cdot m^{\omega})$, and $O(\alpha^{3\omega+1}\cdot m^{\omega})$, respectively.
    As before, we show that for a certain high-range of $\alpha$, the new bound $\alpha^{(j-2)\omega+1}\cdot m^{\omega}$ beats all previous bounds, namely $m^{(j\omega+1)/2}$ and $\alpha^{3j-1}\cdot m$, for $j\geq 2$ (or $\ell=7,10,13,\ldots$). Let $\alpha=\Theta(m^{x})$, where $x\in I_{j}$. We first verify that
    $\alpha^{(j-2)\omega+1}\cdot m^{\omega}\ll m^{(j\omega+1)/2},\text{ or }((j-2)x+1)\omega<j\omega/2,$
    which holds for $x<1/2$. Secondly, we verify that
    $\alpha^{(j-2)\omega+1}\cdot m^{\omega}$ $\ll\alpha^{3j-1}\cdot m,\text{ or }((j-2)x+1)\omega+x<(3j-1)x+1,\text{ or }$
    $x$ $>\frac{\omega-1}{j(3-\omega)+2(\omega-1)},$
    which holds by the choice of the interval $I_{j}$. In particular, for $\ell=7$, this occurs for $x\in\left(\frac{\omega-1}{4},\frac{1}{2}\right)\supset(0.344,0.5)$; for $\ell=10$, this occurs for $x\in\left(\frac{\omega-1}{7-\omega},\frac{1}{2}\right)\supset(0.297,0.5)$; for $\ell=13$, this occurs for $x\in\left(\frac{\omega-1}{10-2\omega},\frac{1}{2}\right)\supset(0.262,0.5)$. Moreover, if $\omega=2$, as conjectured, these intervals extend to $(1/4,1/2)$, $(1/5,1/2)$, and $(1/6,1/2)$, respectively.
  3. (iii)
    $\ell=3j+2$, $j\geq 2$. The refined triangle method [12] with $\ell_{1}=j+1$, $\ell_{2}=j$, $\ell_{3}=j+1$, leads to rectangular matrix multiplication $[O(\alpha^{j-1}m)\times O(\alpha^{j-2}m)]\cdot[O(\alpha^{j-2}m)\times O(\alpha^{j-1}m)]$. Its complexity is at most $O(\alpha^{2})$ times that of the square matrix multiplication with dimension $\alpha^{j-2}m$, the latter of which is $O(\alpha^{(j-2)\omega}\cdot m^{\omega})$. It follows that
    $T(n,m,\ell)=O(\alpha^{2}\cdot\alpha^{(j-2)\omega}\cdot m^{\omega})=O(\alpha^{(j-2)\omega+2}\cdot m^{\omega}).$
    For $\ell=8,11,14$ (the entries in Table 3), these runtimes are $O(\alpha^{2}\cdot m^{\omega})$, $O(\alpha^{\omega+2}\cdot m^{\omega})$, and $O(\alpha^{2\omega+2}\cdot m^{\omega})$, respectively.
    As before, we show that for a certain high-range of $\alpha$, this bound beats all previous bounds, namely $m^{(j\omega+2)/2}$ and $\alpha^{3j}\cdot m$, for $j\geq 2$ (or $\ell=8,11,14,\ldots$). Let $\alpha=\Theta(m^{x})$, where $x\in I_{j}$. We first verify that
    $\alpha^{(j-2)\omega+2}\cdot m^{\omega}\ll m^{(j\omega+2)/2},\text{ or }((j-2)x+1)\omega<j\omega/2,$
    which holds for $x<1/2$. Secondly, we verify that
    $\alpha^{(j-2)\omega+2}\cdot m^{\omega}$ $\ll\alpha^{3j}\cdot m,\text{ or }((j-2)x+1)\omega+2x<3jx+1,\text{or }$
    $x$ $>\frac{\omega-1}{j(3-\omega)+2(\omega-1)},$
    which holds by the choice of the interval $I_{j}$. In particular, for $\ell=8$, this occurs for $x\in\left(\frac{\omega-1}{4},\frac{1}{2}\right)\supset(0.344,0.5)$; for $\ell=11$, this occurs for $x\in\left(\frac{\omega-1}{7-\omega},\frac{1}{2}\right)\supset(0.297,0.5)$; for $\ell=14$, this occurs for $x\in\left(\frac{\omega-1}{10-2\omega},\frac{1}{2}\right)\supset(0.262,0.5)$. Moreover, if $\omega=2$, as conjectured, these intervals extend to $(1/4,1/2)$, $(1/5,1/2)$, and $(1/6,1/2)$, respectively.
    The bound for $\ell=8$, $O(\alpha^{2}\cdot m^{\omega})$, is superseded by the bound $O\left(\alpha^{\omega+1}m^{\frac{\omega+1}{2}}\right)$, using another (so-called “the edge count”) method.
Running time derivations for small $\ell$.
The previous general case derivations together with the instantiations below yield the running times listed in Table 3.
$\ell$ Previous best This paper
$3$ $O(\alpha\cdot m)$ [6], $O(n^{2.372})$ [16], $O(m^{1.407})$ [3]
$4$ $O(\alpha^{2}\cdot m)$ [6], $O(n^{3.334})$ [12], $O(m^{1.686})$ [18] $O(n^{3.251})$
$5$ $O(\alpha^{3}\cdot m)$ [6], $O(n^{4.220})$, $O(m^{2.147})$ [12] $O(n^{4.086})$
$6$ $O(\alpha^{4}\cdot m)$ [6], $O(n^{4.751})$ [12], $O(m^{2.372})$ [12]
$7$ $O(\alpha^{5}\cdot m)$ [6], $O(n^{5.714})$ [12], $O(m^{2.857})$ [12] $O(m^{2.795})$, $O(\alpha\cdot m^{2.372})$
$8$ $O(\alpha^{6}\cdot m)$ [6], $O(m^{3.372})$ [12, 18] $O(m^{3.251})$, $O(\alpha^{3.372}\cdot m^{1.686})$
$9$ $O(\alpha^{7}\cdot m)$ [6], $O(m^{3.558})$ [12, 18] $O(\alpha^{2.372}\cdot m^{2.372})$
$10$ $O(\alpha^{8}\cdot m)$ [6], $O(m^{4.058})$ [12, 18] $O(m^{3.976})$, $O(\alpha^{3.372}\cdot m^{2.372})$
$11$ $O(\alpha^{9}\cdot m)$ [6], $O(m^{4.558})$ [12, 18] $O(m^{4.372})$, $O(\alpha^{4.372}\cdot m^{2.372})$
$12$ $O(\alpha^{10}\cdot m)$ [6], $O(m^{4.744})$ [12, 18] $O(\alpha^{4.744}\cdot m^{2.372})$, $O(\alpha^{6.744}\cdot m^{1.686})$
$13$ $O(\alpha^{11}\cdot m)$ [6], $O(m^{5.160})$ [12, 18] $O(\alpha^{5.744}\cdot m^{2.372})$
$14$ $O(\alpha^{11}\cdot m)$ [6], $O(m^{5.553})$ [12, 18] $O(\alpha^{6.744}\cdot m^{2.372})$
$15$ $O(\alpha^{13}\cdot m)$ [6], $O(m^{5.930})$ [12, 18] $O(\alpha^{7.116}\cdot m^{2.372})$
$16$ $O(\alpha^{14}\cdot m)$ [6], $O(m^{6.338})$ [12, 18] $O(\alpha^{8.115}\cdot m^{2.372})$, $O(\alpha^{10.115}\cdot m^{1.686})$
Table 3: Running time comparison for finding/counting complete subgraphs. The column on new results includes bounds in terms of $m$, based on up-to-date matrix multiplication times, and new arboricity-sensitive bounds in terms of $m$ and $\alpha$; the entries for $\ell=4,5$ in terms of $n$ are also derived in [21].
  1. (i)
    $\ell=3$. The algorithms of Itai and Rodeh [16], and Alon, Yuster, and Zwick [3] for triangle counting/detection run in $O(n^{\omega})$ and $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time, respectively.
  2. (ii)
    $\ell=4$. The refined triangle method [12] with $\ell_{1}=1$, $\ell_{2}=1$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[n\times n]\cdot[n\times m]$ or $[n\times n]\cdot[n\times n^{2}]$ in the worst case. Since $\omega(1,1,2)=\omega(1,2,1)\leq 3.251$, it follows that $T(n,m,4)=O(n^{3.251})$. (See also [22, Table 3]; this entry has been slightly improved in [38].) In particular, this gives a positive answer to a question of Alon, Yuster, and Zwick [3], who asked for an improvement of their $O(n^{\omega+1})=O(n^{3.372})$ upper bound for $K_{4}$ detection. In terms of $m$, by [12, Table 1], $T(n,m,4)=O(m^{1.682})$, but this entry appears unjustified.
  3. (iii)
    $\ell=5$. The refined triangle method [12] with $\ell_{1}=2$, $\ell_{2}=1$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[m\times n]\cdot[n\times m]$ or $[n^{2}\times n]\cdot[n\times n^{2}]$ in the worst case. According to [38, Table 1], $\omega(2,1,2)=2\omega(1,0.5,1)\leq 2\cdot 2.043=4.086$, hence $T(n,m,5)=O(n^{4.086})$.
    On a side note, the extension method [12] with $\ell_{1}=3,\ell_{2}=2$, yields $T(n,m,5)=O(\alpha m\cdot m)=O(\alpha\cdot m^{2})$. Observe however, that $O(\alpha^{3}\cdot m)=O(\alpha\cdot m^{2})$, so the above upper bound does not beat that of the algorithm of Chiba and Nishizeki.
  4. (iv)
    $\ell=6$. The general case derivation yields a runtime of $O(m^{\omega})$.
  5. (v)
    $\ell=7$. The refined triangle method [12] with $\ell_{1}=2$, $\ell_{2}=3$, $\ell_{3}=2$, leads to rectangular matrix multiplication $[m\times\alpha m]\cdot[\alpha m\times m]$ or $[m\times m^{3/2}]\cdot[m^{3/2}\times m]$ in the worst case. Since $\omega(1,1.5,1)\leq 2.795$, it follows that $T(n,m,7)=O(m^{2.795})$.
  6. (vi)
    $\ell=8$. The refined triangle method [12] with $\ell_{1}=\ell_{2}=\ell_{3}=\ell_{4}=2$, leads to counting $K_{4}$’s in a graph with $m$ vertices. Since $\omega(1,1,2)\leq 3.251$, we have $T(n,m,8)=O(m^{3.251})$.
  7. (vii)
    $\ell=9$. The general case derivation yields a runtime of $O(\alpha^{\omega}\cdot m^{\omega})$.
  8. (viii)
    $\ell=10$. By [12, Thm. 2], it takes $O(m^{\beta(10)/2})$, where $\beta(10)=\omega(3,3,4)$. Since according to [22, Table 3], $\omega(1,1.3,1)\leq 2.6217$ and $\omega(1,1.4,1)\leq 2.7084$, by the convexity of $\omega(1,x,1)$ (see [24]), we have $\omega(1,4/3,1)=\omega(1,1.3*2/3+1.4/3,1)\leq 2\omega(1,1.3,1)/3+\omega(1,1.4,1)/3\leq 7.9518/3$. Then $\omega(3,3,4)=3\omega(1,4/3,1)\leq 7.952$, thus $T(n,m,10)=O(m^{\beta(10)/2})=O(m^{3.976})$. Example: for graphs with $\alpha=O(m^{3/8})$, $K_{10}$’s can be counted/detected in $O(m^{3.636})$ time, by the last entry in the right column, whereas the other entries give $O(m^{3.976})$ time or more.
  9. (ix)
    $\ell=11$. By [12, Thm. 2], it takes $O(m^{\beta(11)/2})$, where $\beta(11)=\omega(3,4,4)$. By [38, Table 1], we have $\omega(1,0.75,1)\leq 2.1863$, which yields $\omega(3,4,4)=4\omega(1,0.75,1)\leq 4\times 2.1863\leq 8.746$. Consequently, $T(n,m,11)=O(m^{\beta(11)/2})=O(m^{4.373})$.
  10. (x)
    $\ell=12$. The general case derivation yields a runtime of $O(\alpha^{2\omega}\cdot m^{\omega})$. Example: for graphs with $\alpha=O(m^{0.4})$, $K_{12}$’s can be counted/detected in $O(\alpha^{2\omega}\cdot m^{\omega})=O(m^{1.8\omega})=O(m^{4.269})$ time, whereas the other entries give $O(m^{4.744})$ time or more.
  11. (xi)
    $\ell=13$. By [12, Thm. 2], it takes $O(m^{\beta(13)/2})$, where $\beta(13)=\omega(4,4,5)$. By [22, Table 3], we have $\omega(1,5/4,1)\leq 2.58$, which yields $\omega(4,4,5)=4\omega(1,5/4,1)\leq 4\times 2.58=10.32$. Consequently, $T(n,m,13)=O(m^{\beta(13)/2})=O(m^{5.16})$.
  12. (xii)
    $\ell=14$. By [12, Thm. 2], it takes $O(m^{\beta(14)/2})$, where $\beta(14)=\omega(5,4,5)$. By [38, Table 1], we have $\omega(1,0.8,1)\leq 2.221$, which implies $\omega(4,4,5)=5\omega(1,0.8,1)\leq 5\times 2.221=11.105$, thus $T(n,m,14)=O(m^{\beta(14)/2})=O(m^{5.553})$.
  13. (xiii)
    $\ell=15$. The general case derivation yields a runtime of $O(\alpha^{3\omega}\cdot m^{\omega})$.
  14. (xiv)
    $\ell=16$. By [12, Thm. 2], it takes $O(m^{\beta(16)/2})$, where $\beta(16)=\omega(5,5,6)$. By [38, Table 1], we have $\omega(1,6/5,1)\leq 2.5351$, which implies $\omega(5,5,6)=5\omega(1,6/5,1)\leq 5\times 2.5351\leq 12.676$, thus $T(n,m,16)=O(m^{\beta(16)/2})=O(m^{6.338})$.
The edge count method.
This method combines the extended triangle method with runtime upper bounds of algorithms for the “sparse” case of the detection problem. We illustrate the method for $\ell=4j$, with $j\geq 2$. Detecting the presence of a $K_{\ell}$ amounts to detecting the presence of a $K_{4}$ in a graph $H$ with $O(\alpha^{j-2}m)$ vertices corresponding to the $K_{j}$’s in $G$, where two vertices in $H$ are adjacent if they span $2j$ vertices of $G$ that form a $K_{2j}$. There are $k=O(\alpha^{2j-2}m)$ $K_{2j}$’s in $G$, and so by the result of [18], a $K_{4}$ in $H$ can be detected in $O(k^{(\omega+1)/2})=O\left(\alpha^{(j-1)(\omega+1)}\cdot m^{\frac{\omega+1}{2}}\right)$ time.
Consider three examples: $j=2,3,4$. For $j=2$, the new bound $O\left(\alpha^{\omega+1}\cdot m^{\frac{\omega+1}{2}}\right)$ is better than the earlier bound:. Indeed, it is easily seen that $\alpha^{\omega+1}\cdot m^{\frac{\omega+1}{2}}=O(\alpha^{2}m^{\omega})$ over the entire range of $\alpha$. For $j=3,4$, the new bounds are better than the earlier bounds for $\alpha\ll m^{\frac{\omega-1}{4}}$. Indeed, $\alpha^{2(\omega+1)}\cdot m^{\frac{\omega+1}{2}}\ll\alpha^{2\omega}\cdot m^{\omega}$ and $\alpha^{3(\omega+1)}\cdot m^{\frac{\omega+1}{2}}\ll\alpha^{3\omega+1}\cdot m^{\omega}$ for $\alpha\ll m^{\frac{\omega-1}{4}}$.
These examples are not exhaustive. For instance, one can also apply the edge count method with $\ell=3j$ and use the $O(m^{2\omega/(\omega+1)})=O(m^{1.407})$ time algorithm for triangle detection [3].
5 Conclusion
We conclude by restating a few open problems in the area of finding small complete subgraphs. Recall that triangle detection can be done in $O(n^{\omega})$ time by the algorithm of Itai and Rodeh.
  1. 1.
    Can $K_{3}$ detection be done in $O(n^{2})$ time?
  2. 2.
    Is $K_{3}$ detection as difficult as Boolean matrix multiplication? Initially posed by Alon, Yuster, and Zwick [2], it also appears as in Woeginger’s survey on exact algorithms [39] (as Question 4.3 (c)). V. V. Williams and R. Williams showed that if any of the two problems admits a truly subcubic combinatorial algorithm then also the other one does [37].
  3. 3.
    Can $K_{10}$ detection be accomplished in $O(n^{7.5})$ time? (Question 4.3 (b) in Woeginger’s survey on exact algorithms [39]) We add our own version: Can $K_{100}$ detection be accomplished in $O(n^{75})$ time?
Acknowledgment.
The authors thank an anonymous reviewer for communicating the short folklore proof of Lemma 1.